# Decoding Chess Puzzle Play and Standard Cognitive Tasks for BCI: A Low-Cost EEG Study

### v1

@@ 4.1 Reaction Time and Correctness
Correctness was likewise significantly correlated with workload for all tasks N-Back ($\chi^{2}(3,$ N=6225)=263.89)=263.89) = 263.89, $p<0.001$), Chess ($\chi^{2}(3,$ N=2938)=244.88)=244.88) = 244.88, $p<0.001$), Rotation ($\chi^{2}(3,$ N=1608)=31.50)=31.50) = 31.50, $p<0.001$), and Stroop ($\chi^{2}(1,$ N=4950)=56.20)=56.20) = 56.20, $p<0.001$). However, correctness did not have as strong of an effect across workload within tasks. See Table 4 for post-hoc contrast results. Correctness was significantly different with $p<0.001$ across all contrasts of Chess, for the Stroop contrast, all of N-Back outside of 0-1 ($z=0.97$, $p=0.765$), and for Rotation contrasts 0-3 and 2-3. Rotation contrast of 1-3 was significant ($z=3.09$, $p=0.014$). However, Rotation contrasts of 0-1 ($z=2.12$, $p=0.172$), 0-2 ($z=1.12$, $p=0.754$), and 1-2 ($z=-1.04$, $p=0.765$) were not significant. The results for correctness were as-expected for Chess and Stroop, and largely the case for N-Back. However, given that 0-back and 1-back are effectively trivial tasks, near-perfect accuracy is sensible here. The Rotation task data, however, indicates that the largest degree of rotation (hardest difficulty) was the only one truly distinct from the others in terms of correctness.
Table 4: Post-hoc contrast results for modeling correctness as a function of workload level for each task. Separate models were created for each task type; p-values are corrected using the Benjamini-Hochberg procedure. Outside of N-Back levels 0-1 and differences in Rotation difficulty among 0, 1, and 2, workload level significantly predicts correctness in differences for all tasks.
Task Contrast Est. SE df z p p.adj p.adj.sig
Chess 0 - 3 3.96 0.37 inf 10.76 $<$0.001 $<$0.001 ***
Chess 0 - 2 2.86 0.36 inf 7.87 $<$0.001 $<$0.001 ***
Chess 1 - 3 2.33 0.19 inf 12.55 $<$0.001 $<$0.001 ***
Chess 0 - 1 1.62 0.39 inf 4.19 $<$0.001 $<$0.001 ***
Chess 1 - 2 1.24 0.17 inf 7.15 $<$0.001 $<$0.001 ***
Chess 2 - 3 1.09 0.13 inf 8.69 $<$0.001 $<$0.001 ***
N-Back 0 - 3 2.28 0.18 inf 12.61 $<$0.001 $<$0.001 ***
N-Back 0 - 2 1.63 0.19 inf 8.74 $<$0.001 $<$0.001 ***
N-Back 1 - 3 2.06 0.17 inf 12.19 $<$0.001 $<$0.001 ***
N-Back 0 - 1 0.22 0.23 inf 0.97 0.765 0.765 ns
N-Back 1 - 2 1.41 0.18 inf 8.05 $<$0.001 $<$0.001 ***
N-Back 2 - 3 0.66 0.11 inf 6.02 $<$0.001 $<$0.001 ***
Stroop 0 - 1 1.38 0.18 inf 7.50 $<$0.001 $<$0.001 ***
Rotation 0 - 3 1.38 0.28 inf 4.88 $<$0.001 $<$0.001 ***
Rotation 0 - 2 0.36 0.32 inf 1.12 0.675 0.754 ns
Rotation 1 - 3 0.74 0.24 inf 3.09 0.011 0.014 *
Rotation 0 - 1 0.65 0.30 inf 2.12 0.145 0.172 ns
Rotation 1 - 2 -0.29 0.28 inf -1.04 0.726 0.765 ns
Rotation 2 - 3 1.03 0.26 inf 4.03 $<$0.001 $<$0.001 ***
4.2 Statistical Analysis of Task-Specific Cognitive Load in EEG Signals
Figure 12: EEG power band data as compared across levels of workload within each task. Across workload levels Chess showed significant differences in high-frequency bands ($\beta_{1}$, $\beta_{2}$, $\gamma_{1}$, $\gamma_{2}$). Conversely, N-Back showed significant differences in the lower-frequency bands ($\theta$, $\alpha_{1}$, $\alpha_{2}$). Stroop showed significance in $\alpha_{2}$ and $\beta_{2}$, and Rotation did not show any significant differences across power bands.
Table 5: ANOVA results modeling EEG band data as a function of within-task workload. Separate LMER models were created for each unique combination of wavelength level and task. P-values are corrected with Benjamini-Hochberg procedure across all models. With increased workload levels, Chess showed changes in the upper range of frequency bands, and N-Back showed increased power in the lower range of frequency bands. Only $\beta_{2}$ was the only significant wavelength in the Stroop task, although there is a notably high effect size for $\alpha_{2}$. No changes were observed in the Rotation task.
Task Band NumDF DenDF F p p.adj p.adj.sig $\eta_{p}^{2}$ 95% CI
Chess $\theta$ 3.00 206.78 0.18 0.911 0.911 ns 0.00 [0.00,1]
Chess $\alpha_{1}$ 3.00 207.36 0.35 0.791 0.911 ns 0.01 [0.00,1]
Chess $\alpha_{2}$ 3.00 207.33 2.58 0.055 0.077 ns 0.04 [0.00,1]
Chess $\beta_{1}$ 3.00 204.91 6.27 $<$0.001 0.001 *** 0.08 [0.03,1]
Chess $\beta_{2}$ 3.00 204.27 13.40 $<$0.001 $<$0.001 *** 0.16 [0.09,1]
Chess $\gamma_{1}$ 3.00 205.27 10.29 $<$0.001 $<$0.001 *** 0.13 [0.06,1]
Chess $\gamma_{2}$ 3.00 205.98 8.35 $<$0.001 $<$0.001 *** 0.11 [0.04,1]
N-Back $\theta$ 3.00 141.15 4.90 0.003 0.010 * 0.09 [0.02,1]
N-Back $\alpha_{1}$ 3.00 107.86 6.49 $<$0.001 0.003 ** 0.15 [0.05,1]
N-Back $\alpha_{2}$ 3.00 147.00 4.41 0.005 0.012 * 0.08 [0.02,1]
N-Back $\beta_{1}$ 3.00 112.76 2.13 0.100 0.154 ns 0.05 [0.00,1]
N-Back $\beta_{2}$ 3.00 112.57 1.23 0.303 0.303 ns 0.03 [0.00,1]
N-Back $\gamma_{1}$ 3.00 112.54 1.76 0.159 0.186 ns 0.04 [0.00,1]
N-Back $\gamma_{2}$ 3.00 112.81 2.06 0.110 0.154 ns 0.05 [0.00,1]
Stroop $\theta$ 1.00 39.92 0.00 0.980 0.980 ns 0.00 [0.00,1]
Stroop $\alpha_{1}$ 1.00 34.99 3.15 0.085 0.149 ns 0.08 [0.00,1]
Stroop $\alpha_{2}$ 1.00 34.63 6.58 0.015 0.052 ns 0.16 [0.02,1]
Stroop $\beta_{1}$ 1.00 44.05 0.08 0.772 0.901 ns 0.00 [0.00,1]
Stroop $\beta_{2}$ 1.00 44.52 9.32 0.004 0.027 * 0.17 [0.04,1]
Stroop $\gamma_{1}$ 1.00 40.78 4.43 0.042 0.097 ns 0.10 [0.00,1]
Stroop $\gamma_{2}$ 1.00 44.28 0.17 0.681 0.901 ns 0.00 [0.00,1]
Rotation $\theta$ 3.00 129.00 0.86 0.465 0.926 ns 0.02 [0.00,1]
Rotation $\alpha_{1}$ 3.00 126.15 0.17 0.917 0.926 ns 0.00 [0.00,1]
Rotation $\alpha_{2}$ 3.00 113.15 0.19 0.901 0.926 ns 0.01 [0.00,1]
Rotation $\beta_{1}$ 3.00 130.55 0.26 0.852 0.926 ns 0.01 [0.00,1]
Rotation $\beta_{2}$ 3.00 130.25 1.31 0.272 0.926 ns 0.03 [0.00,1]
Rotation $\gamma_{1}$ 3.00 130.29 0.34 0.796 0.926 ns 0.01 [0.00,1]
Rotation $\gamma_{2}$ 3.00 130.49 0.15 0.926 0.926 ns 0.00 [0.00,1]
4.2.1 Chess
Table 6: Post-hoc contrast results modeling changes in EEG band data as a function of within-task workload for the Chess task. Increases in power were found in among the most separable workload levels (0-3 and 0-2) for all the frequency bands. Increases in power between workload levels 1-3 were likewise found in all except $\beta_{1}$.
Task Band Contrast Est. SE df t p p.adj p.adj.sig $\eta_{p}^{2}$ 95% CI
Chess $\beta_{1}$ 0 - 3 -0.45 0.11 197.61 -3.96 0.001 0.002 ** 0.07 [0.03,1]
Chess $\beta_{1}$ 0 - 2 -0.33 0.11 197.01 -3.01 0.016 0.038 * 0.04 [0.01,1]
Chess $\beta_{1}$ 1 - 3 -0.31 0.11 197.61 -2.72 0.035 0.070 ns 0.04 [0.01,1]
Chess $\beta_{1}$ 0 - 1 -0.14 0.11 197.01 -1.27 0.586 0.669 ns 0.01 [0.00,1]
Chess $\beta_{1}$ 1 - 2 -0.19 0.11 197.01 -1.74 0.305 0.386 ns 0.02 [0.00,1]
Chess $\beta_{1}$ 2 - 3 -0.12 0.11 197.61 -1.02 0.740 0.772 ns 0.01 [0.00,1]
Chess $\beta_{2}$ 0 - 3 -0.72 0.12 197.78 -6.01 $<$0.001 $<$0.001 *** 0.15 [0.08,1]
Chess $\beta_{2}$ 0 - 2 -0.44 0.12 197.01 -3.79 0.001 0.004 ** 0.07 [0.02,1]
Chess $\beta_{2}$ 1 - 3 -0.51 0.12 197.78 -4.26 $<$0.001 0.001 ** 0.08 [0.03,1]
Chess $\beta_{2}$ 0 - 1 -0.21 0.12 197.01 -1.79 0.282 0.376 ns 0.02 [0.00,1]
Chess $\beta_{2}$ 1 - 2 -0.23 0.12 197.01 -2.00 0.193 0.289 ns 0.02 [0.00,1]
Chess $\beta_{2}$ 2 - 3 -0.27 0.12 197.78 -2.30 0.102 0.187 ns 0.03 [0.00,1]
Chess $\gamma_{1}$ 0 - 3 -0.63 0.12 197.80 -5.12 $<$0.001 $<$0.001 *** 0.12 [0.06,1]
Chess $\gamma_{1}$ 0 - 2 -0.39 0.12 197.01 -3.26 0.007 0.019 * 0.05 [0.01,1]
Chess $\gamma_{1}$ 1 - 3 -0.49 0.12 197.80 -4.00 0.001 0.002 ** 0.07 [0.03,1]
Chess $\gamma_{1}$ 0 - 1 -0.14 0.12 197.01 -1.14 0.662 0.722 ns 0.01 [0.00,1]
Chess $\gamma_{1}$ 1 - 2 -0.25 0.12 197.01 -2.12 0.151 0.242 ns 0.02 [0.00,1]
Chess $\gamma_{1}$ 2 - 3 -0.23 0.12 197.80 -1.92 0.223 0.315 ns 0.02 [0.00,1]
Chess $\gamma_{2}$ 0 - 3 -0.56 0.12 197.82 -4.49 $<$0.001 0.001 *** 0.09 [0.04,1]
Chess $\gamma_{2}$ 0 - 2 -0.36 0.12 197.01 -2.95 0.018 0.040 * 0.04 [0.01,1]
Chess $\gamma_{2}$ 1 - 3 -0.46 0.12 197.82 -3.71 0.002 0.005 ** 0.07 [0.02,1]
Chess $\gamma_{2}$ 0 - 1 -0.10 0.12 197.01 -0.80 0.856 0.856 ns 0.00 [0.00,1]
Chess $\gamma_{2}$ 1 - 2 -0.26 0.12 197.01 -2.16 0.140 0.239 ns 0.02 [0.00,1]
Chess $\gamma_{2}$ 2 - 3 -0.20 0.12 197.82 -1.60 0.382 0.459 ns 0.01 [0.00,1]
We observed significant workload effects across all relatively high frequency bands in the Chess task: $\beta_{1}$ ($F_{3,204.91}=6.27$, $p=0.001$, $\eta^{2}_{p}=0.08$), $\beta_{2}$ ($F_{3,204.27}=13.40$, $p<0.001$, $\eta^{2}_{p}=0.16$), $\gamma_{1}$ ($F_{3,205.27}=10.29$, $p<0.001$, $\eta^{2}_{p}=0.13$), and $\gamma_{2}$ ($F_{3,205.98}=8.35$, $p<0.001$, $\eta^{2}_{p}=0.11$). Post-hoc analyses (see Table 6) revealed the most substantial effects in all frequency bands between the largest gaps in workload levels. For the contrast (0-3): $\beta_{1}$: $t_{197.61}=-3.96$, $p=0.002$, $\eta^{2}_{p}=0.07$; $\beta_{2}$: $t_{197.78}=-6.01$, $p<0.001$, $\eta^{2}_{p}=0.15$; $\gamma_{1}$: $t_{197.80}=-5.12$, $p<0.001$, $\eta^{2}_{p}=0.12$; $\gamma_{2}$: $t_{197.82}=-4.49$, $p=0.001$, $\eta^{2}_{p}=0.09$), and the contrast (0-2): $\beta_{1}$: $t_{197.01}=-3.01$, $p=0.038$, $\eta^{2}_{p}=0.04$; $\beta_{2}$: $t_{197.01}=-3.79$, $p=0.004$, $\eta^{2}_{p}=0.07$; $\gamma_{1}$: $t_{197.01}=-3.26$, $p=0.019$, $\eta^{2}_{p}=0.05$; $\gamma_{2}$: $t_{197.01}=-2.95$, $p=0.040$, $\eta^{2}_{p}=0.04$). Further, all except the $\beta_{1}$ band showed significance for the contrasts (1-3): $\beta_{2}$ ($t_{197.78}=-4.26$, $p=0.001$, $\eta^{2}_{p}=0.08$); $\gamma_{1}$ ($t_{197.80}=-4.00$, $p=0.002$, $\eta^{2}_{p}=0.07$); and $\gamma_{2}$ ($t_{197.82}=-3.71$, $p=0.005$, $\eta^{2}_{p}=0.07$).
These findings suggest a strong distinction among the larger separation of workload levels for the Chess task in the high frequency bands. With increased power in the $\beta$ and $\gamma$ frequencies generally associated with active cognitive processing, attention, decision making, and mental effort (Jia and Kohn, 2011; Kropotov, 2016), the progressive increase in high-frequency band power from lower to higher workload conditions supports the interpretation that participants were engaging more cognitive resources as task complexity increased, particularly when comparing across 2 or 3 difficulty levels.
4.2.2 N-Back
Table 7: Post-hoc contrast results modeling EEG band data as a function of within-task workload for the N-Back task. Increases in power were found between workload levels 0-3 for all three bands, and between workload levels and 0-2 for both the $\theta$ and $\alpha_{1}$ bands.
Task Band Contrast Est. SE df t p p.adj p.adj.sig $\eta_{p}^{2}$ 95% CI
N-Back $\theta$ 0 - 3 -0.69 0.22 112.79 -3.09 0.013 0.048 * 0.08 [0.02,1]
N-Back $\theta$ 0 - 2 -0.70 0.22 112.63 -3.20 0.010 0.048 * 0.08 [0.02,1]
N-Back $\theta$ 1 - 3 -0.45 0.23 112.24 -1.97 0.207 0.339 ns 0.03 [0.00,1]
N-Back $\theta$ 0 - 1 -0.24 0.22 113.63 -1.08 0.703 0.953 ns 0.01 [0.00,1]
N-Back $\theta$ 1 - 2 -0.46 0.22 115.15 -2.05 0.177 0.319 ns 0.04 [0.00,1]
N-Back $\theta$ 2 - 3 0.01 0.22 114.33 0.06 1.000 1.000 ns 0.00 [0.00,1]
N-Back $\alpha_{1}$ 0 - 3 -0.67 0.21 111.28 -3.22 0.009 0.048 * 0.09 [0.02,1]
N-Back $\alpha_{1}$ 0 - 2 -0.70 0.20 111.20 -3.42 0.005 0.048 * 0.10 [0.03,1]
N-Back $\alpha_{1}$ 1 - 3 -0.58 0.21 110.28 -2.76 0.034 0.077 ns 0.06 [0.01,1]
N-Back $\alpha_{1}$ 0 - 1 -0.08 0.21 111.94 -0.41 0.977 1.000 ns 0.00 [0.00,1]
N-Back $\alpha_{1}$ 1 - 2 -0.61 0.21 113.12 -2.93 0.021 0.063 ns 0.07 [0.01,1]
N-Back $\alpha_{1}$ 2 - 3 0.03 0.21 112.48 0.15 0.999 1.000 ns 0.00 [0.00,1]
N-Back $\alpha_{2}$ 0 - 3 -0.70 0.22 112.79 -3.14 0.012 0.048 * 0.08 [0.02,1]
N-Back $\alpha_{2}$ 0 - 2 -0.47 0.22 112.64 -2.15 0.143 0.286 ns 0.04 [0.00,1]
N-Back $\alpha_{2}$ 1 - 3 -0.63 0.23 112.25 -2.75 0.034 0.077 ns 0.06 [0.01,1]
N-Back $\alpha_{2}$ 0 - 1 -0.07 0.23 113.64 -0.33 0.988 1.000 ns 0.00 [0.00,1]
N-Back $\alpha_{2}$ 1 - 2 -0.40 0.23 115.16 -1.77 0.291 0.436 ns 0.03 [0.00,1]
N-Back $\alpha_{2}$ 2 - 3 -0.23 0.22 114.33 -1.01 0.741 0.953 ns 0.01 [0.00,1]
In the N-Back task we observed significant workload effects in the lower frequency bands: $\theta$ ($F_{3,141.15}=4.90$, $p=0.010$, $\eta^{2}_{p}=0.09$), $\alpha_{1}$ ($F_{3,107.86}=6.49$, $p=0.003$, $\eta^{2}_{p}=0.15$), and $\alpha_{2}$ ($F_{3,147.00}=4.41$, $p=0.012$, $\eta^{2}_{p}=0.08$). Similar to the differentiations seen with the Chess task, post-hoc analyses (see Table 7) revealed the strongest effect between levels 0-3, represented in the $\theta$ ($t_{112.79}=-3.09$, $p<0.048$, $\eta^{2}_{p}=0.08$), $\alpha_{1}$ ($t_{111.28}=-3.22$, $p<0.048$, $\eta^{2}_{p}=0.09$), and $\alpha_{2}$ ($t_{112.79}=-3.14$, $p=0.048$, $\eta^{2}_{p}=0.08$) frequencies. Effects between workload levels 0-2 were also represented in two out of three bands: $\theta$ ($t_{112.63}=-3.20$, $p=0.048$, $\eta^{2}_{p}=0.08$), and $\alpha_{1}$ ($t_{111.20}=-3.42$, $p=0.048$, $\eta^{2}_{p}=0.10$).
These findings suggest a distinct pattern in the N-Back task with workload effects primarily manifesting in lower frequency bands. This directly confirms other EEG work done with the N-Back task, where an increase in the alpha and theta bands has been found between 0-Back and 2-Back conditions (Brouwer et al., 2012). The prominence of theta and alpha band activity further aligns with established research linking these frequencies to working memory operations, particularly the theta band’s association with repetitive task load (Kropotov, 2009), and the alpha band’s association selective attention, inhibition, and controlled access to stored information (Klimesch, 2012). Similar to the Chess task, the clearest differentiation between workload levels for the N-Back task showed similar effect sizes between the largest workload contrasts, suggesting that the subtlest distinctions in workload are not visible in the data. The overall pattern of increased lower frequency power with heightened working memory load reflects the task’s core demand on memory maintenance and manipulation processes, distinguishing it from the more complex cognitive processing demands observed in the Chess task.
4.2.3 Stroop
Refer to Table 5. In the Stroop task we found significant distinction between congruent and incongruent stimuli (0-1) in the $\beta_{2}$ ($F_{1,44.52}=9.32$, $p=0.027$, $\eta^{2}_{p}=0.17$) band only. However, $\alpha_{2}$ showed near significance with a high effect size ($F_{1,34.63}=6.58$, $p=0.052$, $\eta^{2}_{p}=0.16$); we therefore ran post hoc tests on both. This revealed significance for both, showing $\alpha_{2}$ increasing during the incongruent task ($t_{40.00}=-2.57$, $p<0.014$, $\eta^{2}_{p}=0.14$), and $\beta_{2}$ decreasing during incongruent stimuli ($t_{40.00}=2.57$, $p<0.014$, $\eta^{2}_{p}=0.14$).
Previous work has linked the Stroop task to a decline in beta band power, which may reflect cognitive control mechanisms involved in conflict resolution (Zhao et al., 2015). Supporting this finding, Tafuro et al. (2019) observed a similar decrease in $\beta$ band power during Stroop task performance, suggesting a consistent neural signature of conflict processing. The combination of this observation, along with an increase in $\alpha$ power consistent with the inhibitory mechanisms previously described, suggests a coordinated neural response during conflict resolution in the Stroop task.
4.2.4 Rotation
Notably, our analysis revealed no significant within-task workload differences across difficulty levels in the Rotation task. This consistent pattern of neural engagement suggests that cognitive demands as measured in the prefrontal area may not increase linearly with rotation angles, indicating that participants likely engaged similar neural mechanisms regardless of the rotation magnitude required. It is also possible, however, that the objectivity of the rating scale used may have confounded ‘true’ difficulty. In So et al. (2017)’s similar work, difficulty levels were instead deduced from subjective ratings - it is possible that such difficulty scales might yield better results for this task. Further, analysis of neural signatures from the posterior regions may elicit more differentiable patterns.
4.3 Statistical Analysis of Cross-Task EEG Signal Differentiation
ANOVA results (Table 8) revealed significant differences between tasks across all frequency bands ($F_{3,102.42-106.62}=6.86-26.08$, all $p<0.001$, $\eta^{2}_{p}=0.17-0.43$), with the strongest effects observed in lower frequency ranges: $\alpha_{1}$ ($\eta^{2}_{p}=0.43$), $\alpha_{2}$ ($\eta^{2}_{p}=0.37$), and $\theta$ ($\eta^{2}_{p}=0.36$). Post-hoc analyses (Table 9 and Figure 13) revealed distinct patterns across tasks which are discussed below.

### v2

@@ IV-A Reaction Time and Correctness
Correctness was likewise significantly correlated with workload for all tasks N-Back ($\chi^{2}(3,$ N=6225)=263.89)=263.89) = 263.89, $p<0.001$), Chess ($\chi^{2}(3,$ N=2938)=244.88)=244.88) = 244.88, $p<0.001$), Rotation ($\chi^{2}(3,$ N=1608)=31.50)=31.50) = 31.50, $p<0.001$), and Stroop ($\chi^{2}(1,$ N=4950)=56.20)=56.20) = 56.20, $p<0.001$). However, correctness did not have as strong of an effect across workload within tasks. See LABEL:supp-tab:BEHAVIORAL_CORRECTNESS_POST_HOC in the supplementary material for post-hoc contrast results. Correctness was significantly different with $p<0.001$ across all contrasts of Chess, for the Stroop contrast, all of N-Back outside of 0-1 ($z=0.97$, $p=0.765$), and for Rotation contrasts 0-3 and 2-3. Rotation contrast of 1-3 was significant ($z=3.09$, $p=0.014$). However, Rotation contrasts of 0-1 ($z=2.12$, $p=0.172$), 0-2 ($z=1.12$, $p=0.754$), and 1-2 ($z=-1.04$, $p=0.765$) were not significant. The results for correctness were as-expected for Chess and Stroop, and largely the case for N-Back. However, given that 0-back and 1-back are effectively trivial tasks, near-perfect accuracy is sensible here. The Rotation task data, however, indicates that the largest degree of rotation (hardest difficulty) was the only one truly distinct from the others in terms of correctness.
IV-B Statistical Analysis of Task-Specific Cognitive Load in EEG Signals
Full overall results are shown in the supplementary material in LABEL:supp-tab:WITHIN_EEG_ANOVAS and LABEL:supp-fig:WITHIN_TASK_WORKLOAD.
IV-B1 Chess
We observed significant workload effects across all relatively high frequency bands in the Chess task: $\beta_{1}$ ($F_{3,204.91}=6.27$, $p=0.001$, $\eta^{2}_{p}=0.08$), $\beta_{2}$ ($F_{3,204.27}=13.40$, $p<0.001$, $\eta^{2}_{p}=0.16$), $\gamma_{1}$ ($F_{3,205.27}=10.29$, $p<0.001$, $\eta^{2}_{p}=0.13$), and $\gamma_{2}$ ($F_{3,205.98}=8.35$, $p<0.001$, $\eta^{2}_{p}=0.11$). Post-hoc analyses (see LABEL:supp-tab:WITHIN_EEG_PHOCS_CHESS in the supplementary material) revealed the most substantial effects in all frequency bands between the largest gaps in workload levels. For the contrast (0-3): $\beta_{1}$: $t_{197.61}=-3.96$, $p=0.002$, $\eta^{2}_{p}=0.07$; $\beta_{2}$: $t_{197.78}=-6.01$, $p<0.001$, $\eta^{2}_{p}=0.15$; $\gamma_{1}$: $t_{197.80}=-5.12$, $p<0.001$, $\eta^{2}_{p}=0.12$; $\gamma_{2}$: $t_{197.82}=-4.49$, $p=0.001$, $\eta^{2}_{p}=0.09$), and the contrast (0-2): $\beta_{1}$: $t_{197.01}=-3.01$, $p=0.038$, $\eta^{2}_{p}=0.04$; $\beta_{2}$: $t_{197.01}=-3.79$, $p=0.004$, $\eta^{2}_{p}=0.07$; $\gamma_{1}$: $t_{197.01}=-3.26$, $p=0.019$, $\eta^{2}_{p}=0.05$; $\gamma_{2}$: $t_{197.01}=-2.95$, $p=0.040$, $\eta^{2}_{p}=0.04$). Further, all except the $\beta_{1}$ band showed significance for the contrasts (1-3): $\beta_{2}$ ($t_{197.78}=-4.26$, $p=0.001$, $\eta^{2}_{p}=0.08$); $\gamma_{1}$ ($t_{197.80}=-4.00$, $p=0.002$, $\eta^{2}_{p}=0.07$); and $\gamma_{2}$ ($t_{197.82}=-3.71$, $p=0.005$, $\eta^{2}_{p}=0.07$).
These findings suggest a strong distinction among the larger separation of workload levels for the Chess task in the high frequency bands. With increased power in the $\beta$ and $\gamma$ frequencies generally associated with active cognitive processing, attention, decision making, and mental effort [62, 63], the progressive increase in high-frequency band power from lower to higher workload conditions supports the interpretation that participants were engaging more cognitive resources as task complexity increased, particularly when comparing across 2 or 3 difficulty levels.
IV-B2 N-Back
In the N-Back task we observed significant workload effects in the lower frequency bands: $\theta$ ($F_{3,141.15}=4.90$, $p=0.010$, $\eta^{2}_{p}=0.09$), $\alpha_{1}$ ($F_{3,107.86}=6.49$, $p=0.003$, $\eta^{2}_{p}=0.15$), and $\alpha_{2}$ ($F_{3,147.00}=4.41$, $p=0.012$, $\eta^{2}_{p}=0.08$). Similar to the differentiations seen with the Chess task, post-hoc analyses (see LABEL:supp-tab:WITHIN_EEG_PHOCS_NBACK in the supplementary material) revealed the strongest effect between levels 0-3, represented in the $\theta$ ($t_{112.79}=-3.09$, $p<0.048$, $\eta^{2}_{p}=0.08$), $\alpha_{1}$ ($t_{111.28}=-3.22$, $p<0.048$, $\eta^{2}_{p}=0.09$), and $\alpha_{2}$ ($t_{112.79}=-3.14$, $p=0.048$, $\eta^{2}_{p}=0.08$) frequencies. Effects between workload levels 0-2 were also represented in two out of three bands: $\theta$ ($t_{112.63}=-3.20$, $p=0.048$, $\eta^{2}_{p}=0.08$), and $\alpha_{1}$ ($t_{111.20}=-3.42$, $p=0.048$, $\eta^{2}_{p}=0.10$).
These findings suggest a distinct pattern in the N-Back task with workload effects primarily manifesting in lower frequency bands. This directly confirms other EEG work done with the N-Back task, where an increase in the alpha and theta bands has been found between 0-Back and 2-Back conditions [64]. The prominence of theta and alpha band activity further aligns with established research linking these frequencies to working memory operations, particularly the theta band’s association with repetitive task load [65], and the alpha band’s association selective attention, inhibition, and controlled access to stored information [66]. Similar to the Chess task, the clearest differentiation between workload levels for the N-Back task showed similar effect sizes between the largest workload contrasts, suggesting that the subtlest distinctions in workload are not visible in the data. The overall pattern of increased lower frequency power with heightened working memory load reflects the task’s core demand on memory maintenance and manipulation processes, distinguishing it from the more complex cognitive processing demands observed in the Chess task.
IV-B3 Stroop
Refer to LABEL:supp-tab:WITHIN_EEG_ANOVAS in the supplementary material. In the Stroop task we found significant distinction between congruent and incongruent stimuli (0-1) in the $\beta_{2}$ ($F_{1,44.52}=9.32$, $p=0.027$, $\eta^{2}_{p}=0.17$) band only. However, $\alpha_{2}$ showed near significance with a high effect size ($F_{1,34.63}=6.58$, $p=0.052$, $\eta^{2}_{p}=0.16$); we therefore ran post hoc tests on both. This revealed significance for both, showing $\alpha_{2}$ increasing during the incongruent task ($t_{40.00}=-2.57$, $p<0.014$, $\eta^{2}_{p}=0.14$), and $\beta_{2}$ decreasing during incongruent stimuli ($t_{40.00}=2.57$, $p<0.014$, $\eta^{2}_{p}=0.14$).
Previous work has linked the Stroop task to a decline in beta band power, which may reflect cognitive control mechanisms involved in conflict resolution [67]. Supporting this finding, [68] observed a similar decrease in $\beta$ band power during Stroop task performance, suggesting a consistent neural signature of conflict processing. The combination of this observation, along with an increase in $\alpha$ power consistent with the inhibitory mechanisms previously described, suggests a coordinated neural response during conflict resolution in the Stroop task.
IV-B4 Rotation
Notably, our analysis revealed no significant within-task workload differences across difficulty levels in the Rotation task. This consistent pattern of neural engagement suggests that cognitive demands as measured in the prefrontal area may not increase linearly with rotation angles, indicating that participants likely engaged similar neural mechanisms regardless of the rotation magnitude required. It is also possible, however, that the objectivity of the rating scale used may have confounded ‘true’ difficulty. In [11]’s similar work, difficulty levels were instead deduced from subjective ratings - it is possible that such difficulty scales might yield better results for this task. Further, analysis of neural signatures from the posterior regions may elicit more differentiable patterns.
IV-C Statistical Analysis of Cross-Task EEG Signal Differentiation
ANOVA results (Table LABEL:supp-tab:BETWEEN_EEG_ANOVAS) revealed significant differences between tasks across all frequency bands ($F_{3,102.42-106.62}=6.86-26.08$, all $p<0.001$, $\eta^{2}_{p}=0.17-0.43$), with the strongest effects observed in lower frequency ranges: $\alpha_{1}$ ($\eta^{2}_{p}=0.43$), $\alpha_{2}$ ($\eta^{2}_{p}=0.37$), and $\theta$ ($\eta^{2}_{p}=0.36$). Post-hoc analyses (LABEL:supp-tab:BETWEEN_TASK_PHOCS in the supplementary material and Figure 7) revealed distinct patterns across tasks which are discussed below.